OpenStudy (anonymous):

Complete the ordered pair (0, ? ) so that it is a solution to the linear equation y = 3x + 8 (0, 8) (0, –8/3) (8, 0) (–8/3, 0)

OpenStudy (anonymous):

(0,8)

jimthompson5910 (jim_thompson5910):

The ordered pair (0, ? ) means that x = 0 y = ? (ie y is unknown)

jimthompson5910 (jim_thompson5910):

to find y, we plug in x = 0 and evaluate y = 3x + 8 y = 3*0 + 8 ... replace x with 0 (since x = 0) y = 0 + 8 y = 8

jimthompson5910 (jim_thompson5910):

therefore, the ordered pair goes from (0,?) to (0,8)